### v3

# Strategic Client Selection to Address Non-IIDness in HAPS-enabled FL Networks

Abstract
Federated learning (FL) supported by high-altitude platform stations (HAPS) in non-terrestrial networks (NTN) provides a distinct advantage by allowing the participation of a large number of line-of-sight (LoS) ground clients compared to terrestrial networks, due to their extensive coverage. The presence of many clients enhances the accuracy of the FL model and speeds up convergence. However, the diverse datasets among these clients result in non-independently and identically distributed (non-IID) data, which can negatively affect training accuracy and lead to slower convergence rates. To address this issue, we propose a novel weighted attribute-based client selection strategy that leverages multiple user-specific attributes, including historical traffic patterns, channel qualities, computational capabilities, and prior-round learning performance. By combining these attributes into a composite score for each user at every FL round and selecting users with higher scores as FL clients, the framework ensures more uniform and representative data distributions. Simulation results corroborate the effectiveness of the proposed client selection strategy in enhancing FL model accuracy and convergence rate, as well as reducing training loss.
Index Terms:
Federated learning, HAPS, non-IIDness, client selection.
I Introduction
In the rapidly evolving domain of privacy-preserving machine learning, federated learning (FL) has emerged as a paradigm-shifting approach that develops a global learning model by collaborating among the multitude of clients with unshared data repositories  [1, 2]. The larger the number of repositories, the better FL performs. To achieve this performance, integrating aerial devices from non-terrestrial networks (NTN), such as high altitude platform stations (HAPS), into FL systems has been suggested [3][4]. Because HAPS is located in the stratosphere, it offers a broader geographical reach and direct line-of-sight (LoS) links with ground devices [5]. This capability substantially broadens the set of eligible FL clients by enabling a diverse range of ground devices to participate—an advantage that terrestrial networks typically cannot offer due to their limited coverage and LoS links [6].
However, the large number of clients also introduces the challenge of non-independently and identically distributed (non-IID) data, which arises when each client’s dataset differs substantially from that of other clients, rather than originating from a shared distribution [7]. These differences are often driven by unique traffic usage behaviors, contextual factors, or sensed environmental data, resulting in substantial variability in data distributions across users. Selecting the users with non-IID data distributions to participate in the FL training process degrades the overall effectiveness of the learning, resulting in slower convergence rates and reduced model accuracy [8].
To address this challenge, some recent works [8, 10, 9, 11, 12, 13] have proposed a variety of client selection strategies for FL. Focusing primarily on optimizing operational constraints like efficiency and reliability for FL, these strategies incorporate methods such as client clustering, reputation, game-theoretic, and stochastic techniques. Among these, cluster FL has emerged as a promising solution for addressing non-IID data by grouping clients with similar data distributions [8][9]. For instance, [13] leverages advanced K-means clustering to group clients based on feature similarity, while [9] uses hierarchical clustering for the same purpose. These approaches primarily focus on features derived from the user’s local raw dataset, such as gradients or model updates. However, they often overlook important aspects of the user’s historical network traffic behavior and characteristics, such as past traffic volume and burstiness, which are strongly correlated with their local datasets [14][15].
Figure 1: Network model.
In this work, we present our primary contribution: a weighted attribute-based strategy that selects a more homogeneous set of users by considering features derived from both the user’s local dataset and historical network traffic. This strategy computes a composite score for each user in every communication round by leveraging user-specific attributes, including historical traffic patterns, channel qualities, computational capabilities, and prior-round learning performance. Users with higher composite scores are more likely to possess similar data distributions, making them ideal participants in FL training.
The rationale for selecting these attributes is twofold. First, they help identify users with similar datasets by analyzing historical traffic patterns. This analysis provides insights into traffic usage and users’ prior behavior on the network. Second, the selection focuses on the strongest users among those with similar datasets by assessing their current channel conditions, computational capabilities, and previous-round learning performance, ensuring efficient local training while reducing the impact of stragglers. By consolidating these attributes into a single weighted score, the server can effectively identify the most capable clients with more uniform data, thereby enhancing the performance of FL training.
The remainder of this paper is organized as follows. Section II introduces the network topology and traffic models, and Section III presents the proposed client selection strategy. Section IV discusses the augmented FL model, and Section V provides the simulation results. Finally, Section VI concludes the paper.
II System Model
II-A Network Topology Model
We consider an NTN architecture, as shown in Fig. 1, which consists of a HAPS located in the stratosphere and $K$ ground users [^1] Throughout this work, we use the term “client” to refer to any ground user that has been selected to participate in FL training. that can participate in the FL process. The HAPS acts as an FL server, offering extensive coverage and facilitating communication with the geographically distributed ground users within its footprint. We assume that the locations of these ground users follow a uniform distribution. Each user has local data and is equipped with computational resources, enabling them to undertake local training as part of the FL framework. We assume that HAPS functions as an aerial base station within the proposed NTN architecture, thus having access to historical user network behavior, including user traffic and service demand in terms of traffic volume and burstiness.
II-B Network Traffic Features Modeling
In the following, we discuss the traffic-related features and parameters that form the foundation of our proposed weighted attribute-based strategy.
The packet arrival process can be modeled as a doubly stochastic process, specifically using a Cox (doubly stochastic) Poisson process [16], which accounts for randomness in both the packet arrival rate and the arrival events themselves. Let $N_{k}(t)$ denote the number of packets arriving from user $k$ at the server in a time interval of duration $t$. Under the Cox Poisson process, $N_{k}(t)$ is modeled as a Poisson random variable with a random rate parameter $\lambda_{k}$, and is expressed as
$N_{k}(t)\sim\mathrm{Poisson}\bigl{(}\lambda_{k}\bigr{)},\quad\forall k\in[0,K],\quad t\geq 0.$ (1)
The event rate function $\lambda_{k}$ represents the $k$th user packet arrival rate, which depends on the instantaneous channel realization $h_{k}$ and on the packet size $S_{k}$ (in bits).
Maximum Packet Arrival Rate: Given the packet size $S_{k}$, we define the maximum packet arrival rate (in packets/s) for user $k$ by
$\lambda_{max,k}\;=\;\frac{R_{k}}{S_{k}},$ (2)
where $R_{k}$ is the instantaneous achievable rate of user $k$ (in bits/s) and is computed as
$R_{k}\;=\;b_{k}\,\log_{2}\!\Bigl{(}1+\tfrac{p_{k}|h_{k}|^{2}}{N_{0}b_{k}}\Bigr{)},$ (3)
where $b_{k}$ and $p_{k}$ denote the allocated bandwidth and transmission power, respectively, $N_{0}$ is the noise power spectral density, and $h_{k}$ is the channel coefficient.
Effective Packet Arrival Rate: In practice, the probability of packet loss $P_{\text{loss},k}$ means that the effective rate of packet arrivals is less than $\lambda_{max,k}$. Specifically,
$\lambda_{k}\;=\;\lambda_{max,k}\,\bigl{(}1-P_{\text{loss},k}\bigr{)},$ (4)
where $P_{\text{loss},k}$ is the probability that a packet is lost due to bit errors. For a packet of size $S_{k}$ bits, and assuming independence of bit errors, we can approximate
$P_{\text{loss},k}\;=\;1-\bigl{(}1-P_{b,k}\bigr{)}^{S_{k}}\;\approx\;-\,S_{k}\,\ln\bigl{(}1-P_{b,k}\bigr{)},$ (5)
[…]
Since both the channel power gain $|h_{k}|^{2}$ and the packet size $S_{k}$ are random variables, we derive the average packet arrival rate $\lambda_{k}$ by integrating over their corresponding probability distributions. Formally,
$\mathbb{E}\bigl{[}\lambda_{k}\bigr{]}\;=\;\iint\lambda_{k}\;f\bigl{(}|h_{k}|^{2}\bigr{)}\;f\bigl{(}S_{k}\bigr{)}\;\mathrm{d}|h_{k}|^{2}\,\mathrm{d}s.$ (6)
Under a Rician fading model, the dominant LoS HAPS links cause $|h_{k}|^{2}$ to follow a noncentral chi-square distribution with 2 degrees of freedom, which combines a deterministic LoS component and a stochastic non-LoS (NLoS) component [17]. In a high signal-to-noise ratio (SNR) regime, the deterministic LoS component typically dominates; thus, one can approximate $|h_{k}|^{2}$ by its mean $\Omega_{k}$, where
$\Omega_{k}\;=\;|d_{k}|^{2}+2\,\Psi_{k}^{2},$ (7)
with $|d_{k}|^{2}$ the LoS component’s power, and $\Psi_{k}^{2}$ the NLoS component’s power. Substituting $\Omega_{k}$ into (6) leads to a tractable but approximate expression for $\mathbb{E}[\lambda_{k}]$.
Moreover, empirical studies indicate that packet sizes, modeled as random variables, often follow a log-normal distribution [18]. Consequently, if $S_{k}$ is log-normally distributed with parameters $\mu_{k}$ and $\sigma_{k}^{2}$ [18], then $\mathbb{E}[S_{k}]=e^{\mu_{k}+\tfrac{\sigma_{k}^{2}}{2}}$. Substituting this into the capacity-based arrival rate yields
$\mathbb{E}[\lambda_{k}]\;=\;R_{k}\Bigl{(}e^{-\mu_{k}+\tfrac{\sigma_{k}^{2}}{2}}+\ln\bigl{(}1-P_{b,k}\bigr{)}\Bigr{)},$ (8)
and similarly,
$\mathrm{Var}\bigl{[}\lambda_{k}\bigr{]}\;=\;R_{k}^{2}\,\bigl{(}e^{\sigma_{k}^{2}}-1\bigr{)}\,e^{2(-\mu_{k}+\tfrac{\sigma_{k}^{2}}{2})}.$ (9)
Now, building on the insights discussed above, we define the key traffic-related features as follows:
II-B1 Traffic Volume
We define a key metric, $V_{k}$, as the expected number of packets or traffic volume of user $k$ over an interval $T$. Provided we assume stationarity of the arrival rate $\mathbb{E}[\lambda_{k}]$ within a brief time window $T$, we set
$V_{k}=\mathbb{E}[N_{k}(t)]\;=\;\int_{0}^{T}\mathbb{E}\bigl{[}\lambda_{k}\bigr{]}\;\mathrm{d}t\;=\;T\,\mathbb{E}\bigl{[}\lambda_{k}\bigr{]}.$ (10)
Though real-world network traffic may vary over time, for sufficiently short intervals, the rate can be treated as constant [19], making the analysis more tractable.
II-B2 Traffic Burstiness
Another critical measure is each user’s traffic burstiness, reflecting how aggressively traffic fluctuates relative to its average. Traffic maintaining a steady rate is considered non-bursty, whereas significant peaks and valleys are bursty. We quantify this as
$B_{k}\;=\;\frac{\mathrm{Var}\bigl{[}\lambda_{k}\bigr{]}}{\bigl{(}\mathbb{E}[\lambda_{k}]\bigr{)}^{2}},$ (11)
indicating how large the rate variance is compared to the square of the average rate.
III Proposed Client Selection Strategy
In this section, we discuss each attribute in detail and outline a weighted attributes-based client selection strategy that assigns each user a composite score based on four attributes: traffic metrics, channel quality, computational capability, and dynamic learning performance.
III-A Selection Criteria
Historical Traffic Pattern
To accurately capture users’ historical traffic behavior, we focus on two key metrics: traffic volume $V_{k}$ and burstiness $B_{k}$, both of which can be derived from historical network traffic records available at HAPS. Larger $V_{k}$ typically means higher traffic rates, and higher $B_{k}$ signifies more bursty (less stable) traffic. For client $k$, we combine these into a single metric $\hat{t}$ by first normalizing both $V_{k}$ and $B_{k}$ to $[0,1]$, then defining
$\hat{t}_{k}\;=\;\beta_{V}\,\hat{V}_{k}\;+\;\beta_{B}\,(1-\hat{B}_{k}),\quad\beta_{V}+\beta_{B}=1,$ (12)
where $\hat{V}_{k}$ and $\hat{B}_{k}$ denote the normalized values, and $(\beta_{V},\beta_{B})$ are weighting parameters. A higher $\hat{t}$ thus reflects users with more stable (i.e., less bursty) and consistently high traffic rates (i.e., greater traffic volume), whereas a lower $\hat{t}$ corresponds to users with limited and more volatile traffic patterns. To account for this attribute, $\hat{t}$ is integrated into the final composite score, thereby increasing the likelihood of selecting users with similarly reliable and substantial traffic characteristics for participation in the FL process. Intuitively, users exhibiting similar historical traffic patterns are more likely to possess similar local datasets, thereby reducing data heterogeneity (i.e., less non-IID).
Channel Quality
During the FL process, the channel quality for each client is represented by a normalized measure $\hat{r}_{k}\in[0,1]$, which can be inferred from the available SNR as a function of channel gain at HAPS, i.e., $\frac{p_{k}|h_{k}|^{2}}{N_{0}b_{k}}$. A larger $\hat{r}_{k}$ indicates better channel quality and higher potential throughput. To account for this attribute, $\hat{r}_{k}$ is incorporated into the final composite score to ensure that users with better wireless channel quality are more likely to be selected as FL clients.
Computational Capability
During the FL process, the computing capability $f_{k}$ (in CPU cycles per second) of client $k$ can be determined by HAPS [20]. Let $C_{k}$ denote the number of CPU cycles required to process one data sample. At any communication round, for a given local computation time $l_{k}$ and $J_{k}$ data samples, $f_{k}$ is determined by
$f_{k}=\frac{EC_{k}J_{k}}{l_{k}},\>\forall k,$ (13)
where $E$ is the number of local training epochs. We then normalize $f_{k}$ to obtain $\hat{f}_{k}\in[0,1]$, where a higher $\hat{f}_{k}$ indicates faster processing. Consequently, $\hat{f}_{k}$ is incorporated into the final composite score to ensure that users with greater computing capacity are more likely to be chosen.
Prior-round Learning Performance
During the FL process, HAPS can track each client’s previous-round FL learning performance via a dynamic improvement metric $\hat{m}_{k}$ defined in the following Theorem 1.
Theorem 1.
Let $\Delta_{k}^{(n)}\!\in[0,1]$ be the normalized loss reduction reported by client $k$ after local training round $n$ and choose a memory factor $\zeta\in(0,1)$. Define the recursive update
$\hat{m}_{k}^{(n+1)}\;=\;\zeta\,\hat{m}_{k}^{(n)}+(1-\zeta)\,\Delta_{k}^{(n)}.\vspace{-2mm}$ (14)
Then
  1. (i)
    Optimality. $\hat{m}_{k}^{(n+1)}$ is the unique minimiser of the exponentially weighted least-squares cost $J_{n+1}(\theta)=\sum_{t=0}^{n}\zeta^{\,t}(\theta-\Delta_{k}^{(n-t)})^{2}$.
  2. (ii)
    Closed-form. $\hat{m}_{k}^{(n+1)}=(1-\zeta)\sum_{t=0}^{n}\zeta^{\,t}\Delta_{k}^{(n-t)}+\zeta^{\,n+1}\hat{m}_{k}^{(0)}$, i.e., an exponential moving average (EMA) of past improvements with decay $\zeta$.
  3. (iii)
    Boundedness. If $0\leq\hat{m}_{k}^{(0)}\leq 1$, then $0\leq\hat{m}_{k}^{(n)}\leq 1$ for every $n$.
  4. (iv)
    Convergence. If $\{\Delta_{k}^{(n)}\}$ is mean-stationary with mean $\mu^{\prime}_{k}$, then $\mathbb{E}[\hat{m}_{k}^{(n)}]\to\mu^{\prime}_{k}$ as $n\to\infty$.
Proof.
See Appendix.
Intuitively, a smaller $\zeta$ emphasises the most recent round, whereas a larger $\zeta$ rewards sustained performance.
III-B Composite Score Formulation
After calculating and normalizing all the attributes $\hat{t}_{k}$, $\hat{r}_{k}$, $\hat{f}_{k}$, and $\hat{m}_{k}$ so that higher values are always considered better, HAPS computes a weighted composite score for each attribute using the following linear equation:
$\mathrm{score}_{k}=\varepsilon_{t}\hat{t}_{k}+\varepsilon_{r}\hat{r}_{k}+\varepsilon_{f}\hat{f}_{k}+\varepsilon_{m}\hat{m}_{k},\quad\forall k,$ (15)
This score remains within the range of $[0,1]$ due to the properties of convex combinations. Weights $\varepsilon_{(.)}$ are assigned to each criterion. After each communication round, the absolute Pearson correlation $\beta_{i}$ is calculated for each attribute $i\in\{t,r,f,m\}$ in relation to the reduction in global validation loss on the FL server that was achieved during that round. The corresponding decrease in global validation loss is then recorded as follows:
$y^{(n)}=\frac{\mathcal{L}_{\mathrm{val}}^{(n-1)}-\mathcal{L}_{\mathrm{val}}^{(n)}}{\mathcal{L}_{\mathrm{val}}^{(n-1)}},$ (16)
where $\mathcal{L}_{\mathrm{val}}^{(n)}$ is the global validation loss after the $n$-th aggregation.
Given a memory window of the last $H$ rounds, a paired dataset $\{(x^{(n)}_{i},y^{(n)})\}_{n=n_{0}}^{n_{0}+H-1}$ can be created for every attribute $i$. The relevance score for each attribute is calculated using the absolute Pearson correlation $\beta_{i}=\left|\mathrm{corr}\left(\{x^{(n)}_{i}\}_{n=n_{0}}^{n_{0}+H-1},\{y^{(n)}\}_{n=n_{0}}^{n_{0}+H-1}\right)\right|$. Consequently, the weight vector $\bm{\varepsilon}$ is updated online as
$\varepsilon_{i}=\frac{\beta_{i}}{\sum_{j\in\{t,r,f,m\}}\beta_{j}},\quad i\in\{t,r,f,m\}.$ (17)
This procedure, which operates in $O(|\mathcal{K}|)$, adaptively increases the weight of factors that demonstrate higher predictive power regarding the progress of FL training while ensuring that the constraint $\sum_{i}\varepsilon_{i}=1$ is maintained.
As a result, at round $n$, clients with a composite score exceeding a predefined threshold $\mathrm{score}_{\mathrm{th}}$ are selected
$\mathcal{S}^{(n)}=\{k\mid\mathrm{score}_{k}^{(n)}\geq\mathrm{score}_{\mathrm{th}}\}.$ (18)
By prioritizing clients with higher scores, this selection process favors users whose datasets tend to be more similar. It is important to note that while this approach decreases the number of users participating in the HAPS-aided FL process, the number of selected users remains significantly higher than that of the terrestrial counterpart.
Algorithm 1 Proposed Client Selection Strategy
1:  Initialization: Total number of users $K$, threshold $\text{score}_{\text{th}}$.
2:  for $k=1\to K$ do
3:     Calculate the attribute associated with each selection criterion: $(\hat{t}_{k},\hat{r}_{k},\hat{f}_{k},\hat{m}_{k})$.
4:     Calculate the associated weight vector $\mathbf{\epsilon}=\{\varepsilon_{t},\,\varepsilon_{r},\,\varepsilon_{f},\,\varepsilon_{m}\}$.
5:     Compute composite score for each user using eq. (15).
6:  end for
7:  Select the users with the highest score as FL clients:$\mathcal{S}\;\leftarrow\;\{k|\text{score}_{k}\geq\text{score}_{\text{th}}\}.$
8:  Return Subset $\mathcal{S}$.
IV Augmented FL Model
In this section, we augment the FL algorithm by integrating our proposed novel client selection strategy. Under this algorithm, the global model training proceeds over several communication rounds indexed by $n=0,1,\dots,N-1$. At each round $n$, the algorithm performs the following steps:
  1. 1.
    Client Selection: A subset $\mathcal{S}^{(n)}$ of ground users is selected by HAPS according to our proposed weighted attribute-based selection strategy (presented in Algorithm 1).
  2. 2.
    Model Broadcasting: The HAPS, the FL server, sends the current global model $\mathbf{q}^{(n)}$ to all clients in $\mathcal{S}^{(n)}$.
  3. 3.
    Local Training: Each client $k\in\mathcal{S}^{(n)}$ trains locally on its dataset $\mathcal{D}_{k}$, minimizing
    $\mathcal{L}_{k}\;=\;\mathcal{L}_{\mathrm{CE}}\;+\;\tfrac{\tau}{2}\,\|\mathbf{w}_{k}-\mathbf{q}^{(n)}\|^{2},$ (19)
    where $\tau$ is a constant parameter and $\mathcal{L}_{\mathrm{CE}}$ is the local loss (e.g., cross-entropy). This yields an updated local model $\mathbf{w}_{k}^{(n)}$.
  4. 4.
    Aggregation: The server aggregates the local models $\{\mathbf{w}_{k}^{(n)}:k\in S^{(n)}\}$ to form the new global model as
    $\mathbf{q}^{(n+1)}\;=\;\sum_{k\in S^{(n)}}\alpha_{k}\,\mathbf{w}_{k}^{(n)},$ (20)
    where $\alpha_{k}$ is typically proportional to the local dataset size $\lvert\mathcal{D}_{k}\rvert$.
This sequence of selection, broadcasting, local training, and aggregation repeats until the global model converges to a target accuracy or until a predefined maximum number of rounds $N$ is reached. A concise summary of the algorithmic flow is provided in Algorithm 2.
Algorithm 2 Augmented FL Algorithm
1:  //  Initialization: Training set $\mathcal{D}$, test set $\mathcal{T}$, total clients $K$, total FL rounds $N$, local epochs $E$, learning rate $\eta$, batch size $B$, constant parameter $\tau$, constant $\zeta$.
2:  //  Compute Client Attributes
3:  for $k=1\to K$ do
4:     Derive user traffic features, CSI, and computational capability attributes: $\lambda_{\max,k},\;\lambda_{k},\;P_{\mathrm{loss},k},\;|h_{k}|^{2},f_{k}$.
5:  end for
6:  for $n=0\to N-1$ do
7:     //  Client Selection
[…]
9:     //  Broadcast & Local Training
10:     Send $q^{(n)}$ to each client $k\in S^{(n)}$.
11:     for each client $k\in\mathcal{S}^{(n)}$ in parallel do
12:        Initialize local model $\leftarrow q^{(n)}$.
13:        Train on $\mathcal{D}_{k}$ for $E$ epochs, minimizing: $\mathcal{L}_{k}\;=\;\mathcal{L}_{\mathrm{CE}}\;+\;\frac{\tau}{2}\,\bigl{\|}\mathbf{w}_{k}-\mathbf{q}^{(n)}\bigr{\|}^{2}.$
14:        Obtain local model $\mathbf{w}_{k}^{(n)}$, local loss $\mathcal{L}_{k}$, and improvement metric $\Delta_{k}^{(n)}$.
15:        Set client weight $\alpha_{k}=\frac{|\mathcal{D}_{k}|}{\sum_{j\in S^{(n)}}|\mathcal{D}_{j}|}$.
16:     end for
17:     //  Aggregation
18:     Update the global model: $q^{(n+1)}\;\leftarrow\;\sum_{k\,\in\,S^{(n)}}\alpha_{k}\,\mathbf{w}_{k}^{(n)}.$
19:     //  Dynamic Learning Performance Score Update
20:     for each $k\in S^{(n)}$ do
21:        $m_{k}^{(n+1)}\;\leftarrow\;\zeta\,m_{k}^{(n)}\;+\;\bigl{(}1-\zeta\bigr{)}\,\Delta_{k}^{(n)}.$
22:     end for
23:     //  Evaluation
24:     Evaluate $q^{(n+1)}$ on $\mathcal{T}$, record accuracy $A_{n+1}$, loss $L_{n+1}$, and convergence rate $\gamma_{n+1}=\bigl{|}A_{n+1}-A_{n}\bigr{|}$.
25:  end for
26:  Return Final global model $\,q^{(N)}$ and FL performance metrics $\{A_{n}\}_{n=1}^{N},\{\gamma_{n}\}_{n=1}^{N},\{L_{n}\}_{n=1}^{N}$.
V Simulation Results
In this section, we evaluate the proposed weighted attribute-based client selection strategy using the augmented FL system within a HAPS-aided NTN. The setup involves 500 ground users ($K=500$), each with non-IID data distributions to reflect realistic usage patterns. Specifically, for the non-IID split, we employ a standard label-shard procedure: the 50,000 CIFAR-10 training images are sorted by class, divided into 1,000 equal-sized shards, and two random shards are assigned to each of the 500 clients. As a result, each client’s local dataset contains images from at most two classes. This setup creates a strong, widely used label-skew non-IID scenario that is fully reproducible and comparable with prior FL research.
To emulate real-world conditions, we generate synthetic features such as burstiness and traffic volume, modeled using Rician fading channels and log-normal packet size distributions. This approach captures key variations in user historical traffic behavior. Our primary focus is to investigate how the client selection strategy impacts the outcomes of FL. TABLE I lists the simulation parameters and their values. For a fair comparison, we evaluate our approach against two widely used FL client selection strategies: (i) Random Selection Strategy, which serves as the baseline benchmark commonly adopted in terrestrial FL systems and the idea is to select clients randomly regardless of their resource availability or channel condition, and (ii) Resource-Aware Strategy, which clusters and selects clients based on their available communication resources (e.g., transmission power, bandwidth) and computational resources (e.g., CPU speed) [21]. All figures in this section present a comparative analysis of our proposed strategy against the two benchmark approaches.
In Fig. 2, we examine the average FL test accuracy over communication rounds. As the number of rounds increases, the proposed client selection strategy demonstrates a notable improvement in accuracy, diverging significantly from the benchmark trends. The two benchmark strategies show comparable performance, with the Resource-Aware Strategy offering a slight edge over the Random Selection Strategy. After $n=5$ rounds, our approach outperforms the benchmarks by up to $34\%$, underscoring its effectiveness and high performance.
In Fig. 3, we show the average training loss over the number of FL communication rounds for both our proposed strategy and the benchmarks. The Resource-Aware approach prioritizes clients based on their availability for training, but it does not specifically address non-IID data distributions, limiting its effectiveness in heterogeneous networks. In contrast, our method incorporates a composite score that reflects user traffic behavior, channel conditions, computational capabilities, and historical FL performance, thereby promoting more uniform data distributions among selected clients. As a result, our approach consistently outperforms the benchmarks across all rounds, achieving over $8\%$ improvement in training loss performance.
Finally, in Fig. 4, we examine how quickly the FL model’s performance stabilizes, i.e., the average FL convergence rate, over training communication rounds. Notably, our proposed approach achieves convergence more than $30\%$ faster than both benchmarks after $n=5$ rounds. This improvement highlights the impact of reducing non-IIDness among clients through our client selection strategy, which fosters more homogeneous data distributions and ultimately leads to more efficient and stable training.
TABLE I: Simulation Parameters
Parameter Value
Dataset Non-IID CIFAR-10
Total number of clients $K$ 500
Number of communication rounds $N$ 5
Local training epochs $E$ 2
Batch size 32
Learning rate 0.001
Packet size mean $\mu_{k}$ 7
Packet size standard deviation $\sigma_{k}$ 0.8
Noise power spectral density $N_{0}$ $-174$ dBm/Hz
User transmission power $p_{k}$ $10$ dBm,
Total available bandwidth $20$ MHz
HAPS altitude and broadcasting power $25$ km, $50$ dBm
Score weight $\varepsilon$ 0.25
Score threshold $\text{score}_{\text{th}}$ 0.4
Figure 2: Average FL test accuracy performance over communication rounds.
Figure 3: Average FL training loss performance over communication rounds.
Figure 4: Average FL convergence rate over communication rounds.
VI Conclusion
Our study addresses the challenge of non-IID data in large-scale FL systems by proposing a novel client selection strategy that integrates multiple user-specific attributes, such as traffic patterns, channel conditions, computational capabilities, and historical FL performance, into a weighted composite score. By selecting clients whose data distributions are more uniform, our approach significantly reduces data heterogeneity, resulting in improved training accuracy and training loss, and faster convergence rates. Through extensive experiments and detailed comparisons with baseline methods, we demonstrated the effectiveness and robustness of this strategy, particularly in HAPS-enabled networks.
Proof.
(i) Optimality. Define the normalised, exponentially–weighted least-squares cost
$J_{n+1}(\theta)\;=\;\sum_{t=0}^{n}(1-\zeta)\,\zeta^{t}(\theta-\Delta_{k}^{(n-t)})^{2},\qquad 0<\zeta<1,$ (21)
so that the weights sum to one. Setting $\partial J_{n+1}/\partial\theta=0$ yields $\theta^{\star}=(1-\zeta)\sum_{t=0}^{n}\zeta^{t}\Delta_{k}^{(n-t)},$ which is exactly the EMA update $\hat{m}_{k}^{(n+1)}=\zeta\hat{m}_{k}^{(n)}+(1-\zeta)\Delta_{k}^{(n)}.$
(ii) Closed-form expansion. Substituting $\hat{m}_{k}^{(n)}$ recursively, $\hat{m}_{k}^{(n+1)}=\zeta\bigl{[}\zeta\hat{m}_{k}^{(n-1)}+(1-\zeta)\Delta_{k}^{(n-1)}\bigr{]}+(1-\zeta)\Delta_{k}^{(n)}$ and iterating to $\hat{m}_{k}^{(0)}$ gives
$\hat{m}_{k}^{(n+1)}=(1-\zeta)\sum_{t=0}^{n}\zeta^{t}\Delta_{k}^{(n-t)}+\zeta^{\,n+1}\hat{m}_{k}^{(0)},$ (22)
which is the standard finite-horizon EMA form.
(iii) Boundedness. Assuming the initial value obeys $0\leq\hat{m}_{k}^{(0)}\leq 1$, all weights in the EMA above are non-negative and sum to 1, while every $\Delta_{k}^{(t)}\in[0,1]$. Hence $0\leq\hat{m}_{k}^{(n)}\leq 1$ for every $n$, i.e., the score remains bounded and interpretable.
(iv) Convergence. Let $E_{n}=\mathbb{E}[\hat{m}_{k}^{(n)}]$ and assume the improvement sequence $\{\Delta_{k}^{(n)}\}$ is mean-stationary with $\mu^{\prime}_{k}=\mathbb{E}[\Delta_{k}^{(n)}]$. Taking expectations in the update gives the linear recursion $E_{n+1}=\zeta E_{n}+(1-\zeta)\mu^{\prime}_{k}.$ Its solution is
$E_{n}=\mu^{\prime}_{k}\bigl{(}1-\zeta^{\,n}\bigr{)}+\zeta^{\,n}\hat{m}_{k}^{(0)}\;\xrightarrow[n\to\infty]{}\;\mu^{\prime}_{k},$ (23)
because $10<\zeta<$ 1. Thus, the expected score converges to the long-run mean improvement $\mu^{\prime}_{k}$. ∎